OpenStudy (anonymous):

Use the given conditions to write an equation for the line in point-slope form and general form. Passing through (-8,6) and parallel to the line whose equations is 3x-7y-5=0 The equation of the line in point-slope form is ____

OpenStudy (anonymous):

I got y-6=-3/7(x+8) but the answer is positive 3/7 i dont understand why

OpenStudy (jlastino):

The line is parallel to 3x-7y-5=0 meaning it has the same slope of the given. FInd the slope of 3x-7y=5 -7y=-3x+5 y=3/7x+5 (it's positive 'cause you divided negative by negative)

OpenStudy (anonymous):

omfg lol im slow wow u really have to be meticulous thank you

OpenStudy (jlastino):

np :D